OpenStudy (anonymous):

question on final help !! What is the y-intercept of the line? y = (1/3)x + 2 A. -2 B. -1/3 C. 1/3 D. 2

OpenStudy (anonymous):

to find y intercept, put 0 where you see x y=1/3(0)+2=2

OpenStudy (anonymous):

i got -1/3 ?

OpenStudy (anonymous):

y intercept: x=0 y=1/3(0)+2 y=0+2 y=2